# Gravitational Waves from Massive Black Hole Mergers in ASTRID: Predictions for LISA

### v1

@@ Abstract
We use the ASTRID cosmological simulation to forecast massive black hole binary (MBHB) mergers detectable by LISA. ASTRID offers a detailed approach by directly modeling MBH dynamical friction, allowing for a more realistic tracking of their orbital evolution. Furthermore, it incorporates relatively low-mass MBH seeds down to $5\times 10^{4}$ $\mathrm{M}_{\odot}$, which provides a more complete picture of early MBH formation compared to other cosmological simulations. A key finding is the substantial increase in predicted LISA detection rates when MBH orbital eccentricity is considered, rising from 5.6/yr for circular orbits to 10.5/yr. This enhancement is largely due to the detectability of inspiral sources that will merge after LISA’s observation period, which constitute 46% of detections in the eccentric scenario. This underscores the importance of LISA’s sensitivity to the early inspiral phase, especially for eccentric binaries that emit gravitational waves across a wider frequency band. Most LISA events come from $M_{\mathrm{BH}}\sim 10^{5-6}$ $\mathrm{M}_{\odot}$, low-redshift ($z<2$) and low mass-ratio ($q\sim 0.01-0.1)$ mergers, complementing observations from other instruments like Pulsar Timing Arrays (PTAs). Accounting for eccentricity also broadens the detectable MBH mass range to approximately $10^{9}$ $\mathrm{M}_{\odot}$ from $10^{8}$ $\mathrm{M}_{\odot}$. It shifts the peak of detectable mergers to lower redshifts ($z_{\rm peak}=0.8$ compared to 1.9 for circular orbits). Finally, this study indicates that LISA will detect MBH mergers in a diverse range of host galaxies, including lower-mass galaxies and satellite galaxies. The merger hosts are however fifty times more highly star-forming than the general population for the same mass, and we predict they should form a tight power-law sequence in SFR-$M_{\rm gal}$ space, with SFR= $10^{-8}M_{\rm gal}^{0.9}$ $\mathrm{M}_{\odot}$/yr independent of redshift. This contrasts with expectations for PTA sources, which are more likely to reside in massive central galaxies, implying that LISA will offer unique insights into MBH mergers occurring in different galactic environments. Only $\sim 10\%$ of the remnant MBHs are bright AGNs, and most of them are from massive mergers with $M_{\mathrm{BH}}>10^{6.5}$ $\mathrm{M}_{\odot}$.
@@ 1 Introduction
Massive black holes (MBHs) are ubiquitous in the present Universe, harbored in almost all the massive galactic centers (Kormendy & Richstone, 1995; Magorrian et al., 1998; Kormendy & Ho, 2013). They play a major role in many aspects of galaxy evolution, and have a tight correlation with their host galaxies (Ferrarese & Merritt, 2000; Graham et al., 2001; Ferrarese, 2002; Gültekin et al., 2009; McConnell & Ma, 2013; Reines & Volonteri, 2015; Greene et al., 2016; Schutte et al., 2019). Following the galaxies’ merger, their central MBHs migrate toward the center of the remnant galaxy, losing the orbital energy and momentum via the dynamical friction and three-body scattering. After the MBH orbital decay to sub-parsec scales, the MBH binaries emit gravitational waves (GWs) and eventually coalescence (Merritt, 2013). The GW signal emitted by the merging MBH offers a unique observational pathway for exploring the formation and evolution of MBHs (Wyithe & Loeb, 2003; Sesana et al., 2007; Barausse, 2012).
@@ 2.1 The ASTRID Simulation
MBHs in ASTRID are seeded in halos with $M_{\rm halo,FOF}>5\times 10^{9}~{}h^{-1}\rm{M_{\odot}}$ and $M_{\rm*,FOF}>2\times 10^{6}$ $h^{-1}\rm{M_{\odot}}$, with seed masses stochastically drawn from the range $3\times 10^{4}~{}h^{-1}\rm{M_{\odot}}$ to $3\times 10^{5}~{}h^{-1}\rm{M_{\odot}}$. The gas accretion rate $\dot{M}_{\mathrm{BH}}$ is estimated using the Bondi-Hoyle formalism (Bondi & Hoyle, 1944). Super-Eddington accretion is allowed with an upper limit of twice the Eddington accretion rate. Each MBH radiates with a bolometric luminosity $L_{\rm bol}$ proportional to the accretion rate $\dot{M}_{\mathrm{BH}}$, computed using a mass-to-energy conversion efficiency $\eta=0.1$ (Shakura & Sunyaev, 1973). Of this, 5% of the radiated energy is thermally coupled to the gas within twice the radius of the SPH smoothing kernel of the black hole particle. A subgrid model is applied to estimate MBH dynamical friction following Tremmel et al. (2015) and Chen et al. (2022b). Compared to the more common implementation that directly repositions BHs to the local potential minimum (Weinberger et al., 2017), this model gives well-defined MBH trajectories and velocities. Two MBHs merge if their separation is within two times the gravitational softening length $2e_{g}$, once their kinetic energy is dissipated by dynamical friction and they are gravitationally bound to the local gravitational potential.
@@ 2.2 Shape measurement
In the ASTRID simulation, the trajectories of MBHs are output with high time resolution (every single simulation timestep). Having this fully sampled data leading up to each merger event enables accurate analysis of their orbital evolution. To measure the orbital eccentricity $e$, we employ the shape measurement techniques described in Chen et al. (2022b), which we now briefly summarize. Taking the distance between the two merging black holes $\Delta r$, we define the periapsis $r_{\rm peri}$ to be its local minimum, and the apoapsis $r_{\rm apo}$ to be its local maximum. The eccentricity, $e$, is then given by:
$e=\frac{r_{\rm apo}-r_{\rm peri}}{r_{\rm apo}+r_{\rm peri}}.$ (1)
For each merger, we compute $e$ from the three sets of $r_{\mathrm{apo}}$ and $r_{\mathrm{peri}}$ values that immediately precede the merger. We then select the maximum $e$ from these three values to be the final measurement of $e$ for each merger. This procedure minimizes the impact of falsely identified local extrema, which can produce unrealistically small $e$. Without this choice we would not correctly quantify the effect of incorporating eccentricity when compared to circular orbits. Figure 1 illustrates an example of a highly eccentric merging orbit ($e\approx 1$, upper panel) and a circularized orbit ($e\approx 0$, lower panel) at redshift 0.8. The left panels display $\Delta r$ as a function of time, where the pink and blue crosses indicate the identified $r_{\rm peri}$ and $r_{\rm apo}$ values, which are used for the eccentricity calculation. The right panels show the corresponding orbits projected onto a plane perpendicular to the average relative velocity between the MBHs.
@@ 2.3 Gravitational Wave Signal Calculation
Figure 2: Waveforms for three binaries of different mass and orbital eccentricity $e$ during the 4 years before the coalescence. The strain is calculated for $z=1$, and the mass and the initial $e$ are labeled on the upper right corner. In each panel, the black dotted line is the LISA instrument sensitivity curve with the galactic binary background. The thick black curve shows the strain assuming initial $e=0$. The thin blue curves track the strain at different evolution times. The lines of constant time are shaded from light to dark for early to late evolution. On each blue curve, the circle, square, triangle, and diamond markers represent the evolution of individual harmonics: n=2, 10, 50, and 100, respectively. The thick green curve plots the frequency having the highest strain $f_{\mathrm{peak}}$ for the eccentric waveforms during the evolution.
We use the gwsnrcalc [^1] https://pypi.org/project/gwsnrcalc/ package (Katz & Larson, 2019) to calculated the GW emitted by MBH mergers. We generate two sets of signals, one assuming circular orbits (labeled as Circ), and the other includes the orbital eccentricity measured in Section 2.2 (labeled as Ecc).
@@
For Ecc, gravitational wave strain is influenced by the orbital eccentricity. The $h_{\mathrm{s}}$ from an individual eccentric source can be related to that of a circular sources as (see Amaro-Seoane et al., 2010; Kelley et al., 2017; Chen et al., 2022b):
@@
PhenomD describes the waveforms in the inspiral-merger-ringdown (IMR) phase in the $l=m=2$ quasinormal mode. For an eccentric binary merger, the inspiral waveforms are sufficiently accurate, but waveform models including the contributions from the merger and ringdown stages are less developed. Although recent efforts have been made to construct IMR waveforms with eccentricity (Hinder et al., 2018; Setyawati & Ohme, 2021; Chattaraj et al., 2022), some important physical effects such as spin or higher order modes are not included, and there is still a lack of a clear and simple model to be implemented. Hence, to make a better comparison between the Circ and Ecc runs, in this work we only consider the waveforms from the inspiral phase while not including the emission during the phase of merger and ringdown.
@@
In Figure 2, we show the waveforms for three binaries with different masses and orbital eccentricities $e$ during the four years before coalescence. The strain is calculated for $z=1$, and the mass and the initial $e$ are labeled in the upper right corner. In each panel, the black dotted line is the LISA instrument sensitivity curve with the galactic binary background. The thick black curve shows the strain assuming initial $e=0$. The thin blue curves track the strain at different evolution times. The lines of constant time are shaded from light to dark for early to late evolution. On each blue curve, the circle, square, triangle, and diamond markers represent the evolution of individual harmonics: n=2, 10, 50, and 100, respectively. The thick green curve plots the frequency having the highest strain $f_{\mathrm{peak}}$ for the eccentric waveforms during the evolution. For the binaries with the circular orbit, the strain comes from the individual harmonic mode $n=2$; while for the eccentric binaries, the GW signal is contributed by all $n\geq 1$ harmonic modes, which produces a wide frequency range. As we will discuss later in Section 2.4, this boosts the detectability for some massive or low-$z$ sources. The eccentricity decays with the GW emission (Peters & Mathews, 1963), which is called circularization. This makes the waveforms shortly before the coalescence similar to those from circular binaries (black thick curves), and the dominant frequency $f_{\mathrm{peak}}$ evolves towards the $n=2$ mode (blue circles).
@@ 2.4 Merger rate prediction
An advantage of implementing this Monte Carlo analysis is that we can include the GW signals emitted by the inspiral sources, which coalesce after the LISA mission ends. To do this, we sample the MBH populations merging within $t_{\mathrm{dur}}$, and the first $t_{\mathrm{obs}}$ phase $(t_{\mathrm{dur}}>t_{\rm obs})$ corresponds to the observation time. The binaries which take less time than $t_{\mathrm{obs}}$ to merge are marked as ‘merger sources’ (these are observed to coalesce). The binaries which merge after $t_{\mathrm{obs}}$ are labeled as ‘inspiral sources’ as they keep inspiraling during the whole observation phase. For the LISA mission, $t_{\mathrm{obs}}=4$ years. Following Katz et al. (2020), we set $t_{\mathrm{dur}}=100$ years  [^2] We test different choices of $t_{\mathrm{dur}}$, and confirm that our prediction results converge with $t_{\mathrm{dur}}\geq 100$ yr.. As the Universe evolves negligibly over $t_{\mathrm{dur}}$, we can assume that the merger times of the $N_{\mathrm{t_{\rm dur}}}$ sources are uniformly distributed over the interval of length $t_{\mathrm{dur}}$. We therefore draw the observation start $t_{\mathrm{start}}$ randomly from between 00 and $100$ years to decide which 4-year phase for an individual merger is observed during $t_{\mathrm{dur}}$ beyond coalescence, For the observation end time $t_{\mathrm{end}}$, $t_{\mathrm{end}}=t_{\mathrm{start}}-t_{\mathrm{obs}}$ if $t_{\mathrm{start}}>t_{\mathrm{obs}}$ (inspiral sources); and $t_{\rm end}=0$ if $t_{\mathrm{start}}\leq t_{\mathrm{obs}}$ (merger sources). GW signals from $t_{\rm start}$ to $t_{\rm end}$ before the coalescence are used to calculate the SNR for each merger.
Figure 3: Illustration of ASTRID at $z=0.8$. The upper panel shows the large-scale structure in a slice of $100\times 60\times 10\ \mathrm{Mpc}^{3}\,h^{-3}$, with gas temperature represented by color: hotter regions appear in red, while cooler regions are shown in blue. Yellow circles mark the locations of detectable mergers (i.e., with SNR $>10$) occurring within the 670 Myr around $z=0.8$. Five merging systems are selected as examples, whose positions are marked in the plot. The middle panels show the stellar density field within $600\ \mathrm{Mpc}/h$ around the merging MBH binaries, color-coded by the stellar age, where blue regions represent young stars. The red crosses show the position of primary MBH (or the remnant MBH), the green crosses show the secondary MBH. The bottom panels show the host galaxy of the merging MBHs, with the RGB channels representing the flux in the rest-frame $grz$ color bands.
@@
Figure 5: Probability density function (PDF) of merger orbit eccentricities $e$ measured based on Eq. 1. Left: PDF for mergers at different redshifts. Solid curves represent the distribution from ASTRID merger population, and dotted curves show the fitted $\beta$ distribution (Equ. 9) for the redshifts with the corresponding color. Right: PDF for mergers with different total MBH masses.
Figure 6: Strain and frequency distribution for the mergers in ASTRID. The left frame is for Circ run and the right panel is for Ecc run. For each MBH merger, the plotted frequency and strain correspond to the evolution time point and harmonic mode that makes the largest contribution to SNR within 4 years before the coalescence. For the Circ run, the harmonic mode is always $n=2$. We color the distribution based on the redshift and the number of events, as shown by the right colorbar: the high(low) redshifts are represented by the blue(red) colors, and the darkness corresponds to the regions with very few events. The redshift is averaged over the merger population in each pixel.
Figure 7: Parameter space of MBH mass and redshift explored by LISA. The left panel plots the two-dimensional distribution of the redshift and the total mass of merging MBHs $M_{\mathrm{tot}}$. The right panel shows the distribution of the mass of the primary MBH $M_{1}$ and the secondary MBH $M_{2}$. The underlying blue color gradient represents the differential number of detected mergers during the LISA 4-year observation for Ecc run: $dN/(d\log M\,dz\,dt)\times\mathrm{4yr}$ for the left panel and $dN/(d\log M_{1}\,d\log M_{2}\,dt)\times\mathrm{4yr}$ for the right panel. The black and red contours correspond to the detected MBH merger population for Ecc run and Circ run, respectively.

### v2

@@ Abstract
We use the ASTRID cosmological simulation to forecast massive black hole (MBH) mergers detectable by LISA down to $z=0$. ASTRID directly models MBH dynamical friction, allowing a realistic tracking of their trajectory. It also incorporates relatively low-mass MBH seeds down to $5\times 10^{4}$ $\mathrm{M}_{\odot}$, providing a more complete picture of LISA MBH mergers. We find that LISA MBH mergers initially have high eccentricities, peaking around $e_{0}=0.8$ across all redshifts. Accounting for this boosts the event rate from 5.6/yr (if circular orbits are assumed) to 10.5/yr. This enhancement is largely due to additional inspiral sources that will coalesce after LISA’s observation, which constitute 46% of detected events. This underscores the importance of LISA’s sensitivity to the early inspiral phase, especially for eccentric binaries that emit gravitational waves across a wider frequency band. Most LISA events in ASTRID arise from $M_{\mathrm{BH}}\sim 10^{5-6}$ $\mathrm{M}_{\odot}$, low-redshift ($z<2$) and low mass-ratio ($q\sim 0.01-0.1)$ mergers. Accounting for eccentricity broadens the detectable MBH mass range up to $10^{9}$ $\mathrm{M}_{\odot}$  and shifts the peak of detectable mergers to a lower redshift $z_{\rm peak}=0.8$. This implies that the most massive LISA events may also be PTA sources. We predict LISA events to be in various galaxy environments, including many low-mass satellite galaxies. The EM counterparts of most LISA sources have AGN luminosities $L_{\rm bol}>10^{42}$ erg/s, albeit only $1\%$ with $>10^{44}$ erg/s. The brightest AGN are those associated with the rare LISA/PTA events with $M_{\rm BH}>10^{8}$ $\mathrm{M}_{\odot}$.
@@ 1 Introduction
Massive black holes (MBHs) are ubiquitous in the present Universe, harbored in almost all the massive galactic centers (Kormendy & Richstone, 1995; Magorrian et al., 1998; Kormendy & Ho, 2013). They play a major role in many aspects of galaxy evolution, and have a tight correlation with their host galaxies (Ferrarese & Merritt, 2000; Graham et al., 2001; Ferrarese, 2002; Gültekin et al., 2009; McConnell & Ma, 2013; Reines & Volonteri, 2015; Greene et al., 2016; Schutte et al., 2019). Following the galaxies’ merger, their central MBHs migrate toward the center of the remnant galaxy, losing the orbital energy and momentum via dynamical friction (DF) and three-body scattering. After the MBH orbital decay to sub-parsec scales, the MBH binaries emit gravitational waves (GWs) and eventually coalescence (Merritt, 2013). The GW signal emitted by the merging MBH offers a unique observational pathway for exploring the formation and evolution of MBHs (Wyithe & Loeb, 2003; Sesana et al., 2007; Barausse, 2012).
@@ 2.1 The ASTRID Simulation
MBHs in ASTRID are seeded in halos with $M_{\rm halo,FOF}>5\times 10^{9}~{}h^{-1}\rm{M_{\odot}}$ and $M_{\rm*,FOF}>2\times 10^{6}$ $h^{-1}\rm{M_{\odot}}$, with seed masses stochastically drawn from the range $3\times 10^{4}~{}h^{-1}\rm{M_{\odot}}$ to $3\times 10^{5}~{}h^{-1}\rm{M_{\odot}}$. The gas accretion rate $\dot{M}_{\mathrm{BH}}$ is estimated using the Bondi-Hoyle formalism (Bondi & Hoyle, 1944). Super-Eddington accretion is allowed with an upper limit of twice the Eddington accretion rate. Each MBH radiates with a bolometric luminosity $L_{\rm bol}$ proportional to the accretion rate $\dot{M}_{\mathrm{BH}}$, computed using a mass-to-energy conversion efficiency $\eta=0.1$ (Shakura & Sunyaev, 1973). Of this, 5% of the radiated energy is thermally coupled to the gas within twice the radius of the SPH smoothing kernel of the black hole particle. A subgrid model is applied to estimate MBH dynamical friction following Tremmel et al. (2015) and Chen et al. (2022b). Compared to the more common implementation that directly repositions BHs to the local potential minimum (Weinberger et al., 2017), this model gives well-defined MBH trajectories and velocities. Two MBHs merge if their separation is within two times the gravitational softening length $2\epsilon_{g}$, once their kinetic energy is dissipated by DF and they are gravitationally bound to the local gravitational potential.
@@ 2.2 Shape Measurement
In the ASTRID simulation, the trajectories of MBHs are output with high time resolution (every single simulation timestep). Having this fully sampled data leading up to each merger event enables accurate analysis of their orbital evolution. To measure the orbital eccentricity $e_{0}$, we employ the shape measurement techniques described in Chen et al. (2022b), which we now briefly summarize. We use the subscript of 00 to indicate the orbital eccentricity is used as the initial condition in the evolution during the GW regime, which will be changed due to the GW circularization (Peters & Mathews, 1963). Taking the distance between the two merging black holes $\Delta r$, we define the periapsis $r_{\rm peri}$ to be its local minimum, and the apoapsis $r_{\rm apo}$ to be its local maximum. The eccentricity, $e_{0}$, is then given by:
$e_{0}=\frac{r_{\rm apo}-r_{\rm peri}}{r_{\rm apo}+r_{\rm peri}}.$ (1)
For each merger, we compute $e_{0}$ from the three sets of $r_{\mathrm{apo}}$ and $r_{\mathrm{peri}}$ values that immediately precede the merger. We then select the maximum $e_{0}$ from these three values to be the final measurement of $e_{0}$ for each merger. This procedure minimizes the impact of falsely identified local extrema, which can produce unrealistically small $e_{0}$. Without this choice, we would not correctly quantify the effect of incorporating eccentricity when compared to circular orbits. Figure 1 illustrates an example of a highly eccentric merging orbit ($e_{0}\approx 1$, upper panel) and a circularized orbit ($e_{0}\approx 0$, lower panel) at redshift 0.8. The left panels display $\Delta r$ as a function of time, where the pink and blue crosses indicate the identified $r_{\rm peri}$ and $r_{\rm apo}$ values, which are used for the eccentricity calculation. The right panels show the corresponding orbits projected onto a plane perpendicular to the average relative velocity between the MBHs.
@@ 2.3 Gravitational Wave Signal Calculation
Figure 2: Waveforms for three binaries of different mass and orbital eccentricity $e_{0}$ during the 4 years before the coalescence. The strain is calculated for $z=1$, and the mass and the initial $e_{0}$ are labeled on the upper right corner. In each panel, the black dotted line is the LISA instrument sensitivity curve with the galactic binary background. The thick black curve shows the strain assuming initial $e_{0}=0$. The thin blue curves track the strain at different evolution times. The lines of constant time are shaded from light to dark for early to late evolution. On each blue curve, the circle, square, triangle, and diamond markers represent the evolution of individual harmonics: n=2, 10, 50, and 100, respectively. The thick green curve plots the frequency having the highest strain $f_{\mathrm{peak}}$ for the eccentric waveforms during the evolution.
We use the gwsnrcalc [^1] https://pypi.org/project/gwsnrcalc/ package (Katz & Larson, 2019) to calculate the GW emitted by MBH mergers. We generate two sets of GW signals, one assuming circular orbits (labeled as Circ), and our main prediction, which incorporates the orbital eccentricity inferred (and evolved) from the BH orbits in ASTRID, as described in Section 2.2. We label the predictions using the eccentricity information as Ecc.
@@
For Ecc, gravitational wave strain is influenced by the orbital eccentricity. The $h_{\mathrm{s}}$ from an individual eccentric source can be related to that of a circular source as (see Amaro-Seoane et al., 2010; Kelley et al., 2017; Chen et al., 2022b):
@@
PhenomD describes the waveforms in the inspiral-merger-ringdown (IMR) phase in the $l=m=2$ quasinormal mode. For an eccentric binary merger, the inspiral waveforms are sufficiently accurate, but waveform models including the contributions from the merger and ringdown stages are less developed. Although recent efforts have been made to construct IMR waveforms with eccentricity (Hinder et al., 2018; Setyawati & Ohme, 2021; Chattaraj et al., 2022), some important physical effects such as spin or higher order modes are not included, and there is still a lack of a clear and simple model to be implemented. Hence, to make a better comparison between the Circ and Ecc, in this work we only consider the waveforms from the inspiral phase while not including the emission during the phase of merger and ringdown.
@@
In Fig. 2, we show the waveforms for three binaries with different masses and orbital eccentricities $e_{0}$ during the four years before coalescence. The strain is calculated for $z=1$, and the mass and the initial $e_{0}$ are labeled in the upper right corner. In each panel, the black dotted line is the LISA instrument sensitivity curve with the galactic binary background. The thick black curve shows the strain assuming an initial $e=0$. The thin blue curves track the strain at different evolution times. The lines of constant time are shaded from light to dark for early to late evolution. On each blue curve, the circle, square, triangle, and diamond markers represent the evolution of individual harmonics: n=2, 10, 50, and 100, respectively. The thick green curve plots the frequency having the highest strain $f_{\mathrm{peak}}$ for the eccentric waveforms during the evolution. For the binaries with the circular orbit, the strain comes from the individual harmonic mode $n=2$; while for the eccentric binaries, the GW signal is contributed by all $n\geq 1$ harmonic modes, which produces a wide frequency range. If the circularization causes the eccentricity close to zero in the end (the left and middle panels), we find that the behavior of the Ecc is very similar to that of the Circ. Since the total SNR is typically dominated by the emission in the short phase before coalescence, these fully circularized sources have similar detectability in both the Ecc and Circ. While for the massive mergers shown in the right panel, incorporating eccentricity boots the detectability. This is because the high emission modes produce emissions closer to the high-sensitivity band, contributing significantly to the total SNR.
@@ 2.4 Merger rate prediction
An advantage of implementing this Monte Carlo analysis is that we can include the GW signals emitted by the inspiral sources, which coalesce after the LISA mission ends. To do this, we sample the MBH populations merging within $t_{\mathrm{dur}}$, and the first $t_{\mathrm{obs}}$ phase $(t_{\mathrm{dur}}>t_{\rm obs})$ corresponds to the observation time. The binaries which take less time than $t_{\mathrm{obs}}$ to merge are marked as ‘merger sources’, which are observed to coalesce. The binaries which merge after $t_{\mathrm{obs}}$ are labeled as ‘inspiral sources’ as they keep inspiraling during the whole observation phase. For the LISA mission, $t_{\mathrm{obs}}=4$ years. Following Katz et al. (2020), we set $t_{\mathrm{dur}}=100$ years  [^2] We test different choices of $t_{\mathrm{dur}}$, and confirm that our prediction results converge with $t_{\mathrm{dur}}\geq 100$ yr.. As the Universe evolves negligibly over $t_{\mathrm{dur}}$, we can assume that the merger times of the $N_{\mathrm{t_{\rm dur}}}$ sources are uniformly distributed over the interval of length $t_{\mathrm{dur}}$. We therefore draw the observation start $t_{\mathrm{start}}$ randomly from between 00 and $100$ years to decide which 4-year phase for an individual merger is observed during $t_{\mathrm{dur}}$ beyond coalescence, For the observation end time $t_{\mathrm{end}}$, $t_{\mathrm{end}}=t_{\mathrm{start}}-t_{\mathrm{obs}}$ if $t_{\mathrm{start}}>t_{\mathrm{obs}}$ (inspiral sources); and $t_{\rm end}=0$ if $t_{\mathrm{start}}\leq t_{\mathrm{obs}}$ (merger sources). GW signals from $t_{\rm start}$ to $t_{\rm end}$ before the coalescence are used to calculate the SNR for each merger. We use the $e_{0}$ measured by equation 1 as the initial eccentricity at the beginning of LISA observation ($t_{\rm start}$), and it will be circularized with the GW emission according to equation 5.13 in Peters (1964). We do not include the effect of circularization before the observation starts as the eccentricity decay rate, which depends on $a^{-4}$ ($a$ is the length of semi major axis of the binary orbit), is typically small during the early inspiral phase.
@@
Figure 4: Illustration of ASTRID at $z=0.8$. The upper panel shows the large-scale structure in a slice of $100\times 60\times 10\ \mathrm{Mpc}^{3}\,h^{-3}$, with gas temperature represented by color: hotter regions appear in red, while cooler regions are shown in blue. Yellow circles mark the locations of detectable mergers (i.e., with SNR $>10$) occurring within the 670 Myr around $z=0.8$. Five merging systems are selected as examples, whose positions are marked in the plot. The middle panels show the stellar density field within $600\ \mathrm{kpc}/h$ around the merging MBH binaries, color-coded by the stellar age, where blue regions represent young stars. The red crosses show the position of the primary MBH (or the remnant MBH), the green crosses show the secondary MBH. The bottom panels show the host galaxy of the merging MBHs, with the RGB channels representing the flux in the rest-frame $grz$ color bands.
Figure 5: Probability density function (PDF) of merger orbit eccentricities $e_{0}$ measured based on Eq. 1. Left: PDF for mergers at different redshifts. Solid curves represent the distribution from ASTRID merger population, and dotted curves show the fitted $\beta$ distribution (Equ. 9) for the redshifts with the corresponding color. Right: PDF for mergers with different total MBH masses.
Figure 6: Strain and frequency distribution for the mergers in ASTRID. The left frame is for Circ and the right panel is for Ecc. For each MBH merger, the plotted frequency and strain correspond to the evolution time point and harmonic mode that makes the largest contribution to SNR within 4 years before the coalescence. For the Circ, the harmonic mode is always $n=2$. We color the distribution based on the redshift and the number of events, as shown by the right colorbar: the high(low) redshifts are represented by the blue(red) colors, and the darkness corresponds to the regions with very few events. The redshift is averaged over the merger population in each pixel.
Figure 7: Parameter space of MBH mass and redshift explored by LISA. The left panel plots the two-dimensional distribution of the redshift and the total mass of merging MBHs $M_{\mathrm{tot}}$. The right panel shows the distribution of the mass of the primary MBH $M_{1}$ and the secondary MBH $M_{2}$. The underlying blue color gradient represents the differential number of detected mergers during the LISA 4-year observation for Ecc: $dN/(d\log M\,dz\,dt)\times\mathrm{4yr}$ for the left panel and $dN/(d\log M_{1}\,d\log M_{2}\,dt)\times\mathrm{4yr}$ for the right panel. The black and red contours correspond to the detected MBH merger population for Ecc and Circ, respectively.